CAVITY RINGDOWN SPECTROSCOPY OF HC_{7}H AND HC_{9}H

\underline{C. D. BALL}, M. C. MCCARTHY, and P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, Cambridge, MA 02138 and Division of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138.

Vibronic bands of the^{3}Σ_{u}^{−} ← X^{3}Σ_{g}^{−} transitions of linear HC_{7}H and HC_{9}H have been detected in the gas phase using cavity ringdown spectroscopy in the wavelength range 500-600 nm. The 0_{0}^{0} bands of both molecules, as well as an additional vibronic band of HC_{9}H, were measured. Both molecules were produced in a DC discharge through a dilute mixture of diacetylene in Ar in the throat of a supersonic nozzle. The identifications were confirmed by detecting the singly and doubly deuterated species of both molecules, and by comparison with previous matrix data^{a}. Neither molecule could be detected by laser-induced fluorescence. Each band exhibits expected P- and R-branches, and the band origins have been measured to 0.1 cm^{−1} or better. For HC_{9}H, the wavelength of the 0_{0}^{0} band agrees to within one part in 10^{4} with that of the weak diffuse interstellar band at 581.86 nm. Astronomical verification of this assignment could be achieved by detecting the expected P- and R-branches in the DIB, or by detecting the somewhat weaker ν_{3} band at 521.91 nm.

aJ. Fulara, P. Freivogel, D. Forney, and J. P. Maier, J. Chem. Phys. 103, 8805 (1995).